In garden peas axial flowers (A) are dominant over terminal flower (a). Two plants each heterozygous for this gene are crossed. What is the probability that first four offspring from this cross will have axial flowers?
Answer #1

Answer: 81/256 = 31.6%
Explanation:
Axial = Dominant = A
Terminal = Recessive = a
Parental cross: Aa X Aa
Progeny: AA Aa Aa aa
The probability for axial flowers in the progeny = 3/4
The probability that the first four offspring from this cross will have axial flowers = 3/4 X 3/4 X 3/4 X 3/4
= 81/256
= 0.316
= 31.6%

Please rate the solution. Your efforts are highly appreciated. Thank you.